. Sum of two numbers is 95. If one exceeds the other by 15, find the numbers. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let one of the numbers be= x.
Then, the other number becomes x + 15 According to the question, x+x+15=95
⇒2x+15=95
⇒2x=95-15
⇒2x=80
⇒x=80/2
⇒x=40
First number = x = 40
And, other number = x + 15 = 40 + 15 = 55

Thanks